OpenStudy (anonymous):

use the quotient rule with Generalized power rule to find the exact value of f'(2) for the function f'(x)=(4x+3)^3/(2x-1)

OpenStudy (anonymous):

just because I am curious, f(x) is a function, or is it already a derivative? Wether or not, it will change nothing about the answer. \[ \large f'(x)= \frac{3(4x+3)^2\cdot 4(2x-1) -2\cdot (4x+3)^3}{(2x-1)^2} \]

OpenStudy (anonymous):

Just here again for reminding, the quotient rule is: \[ \left(\frac{u}{v}\right)'=\frac{u'v-uv'}{v^2} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but I thought (2x-1) suspost to came first on top.

OpenStudy (anonymous):

\[ u= (4x+3)^3 \rightarrow \frac{du}{dx}=u'=3(4x+3)^2 \cdot 4 = 12(4x+3)^2 \\v=(2x-1) \rightarrow v'=2 \]

OpenStudy (anonymous):

the order of multiplication doesn't matter if that's what you mean, just be careful with subtraction

OpenStudy (anonymous):

still don't get it what did you get for your anserw?

OpenStudy (anonymous):

\[ \frac{12(4x+3)^2(2x-1)-2(4x+3)^3}{(2x-1)^2}\]

OpenStudy (anonymous):

Do you know how to apply the chain rule in here?

OpenStudy (anonymous):

it just says use quotient rule and generalized power rule to find the exact value.

OpenStudy (jiteshmeghwal9):

do u mean factor theorem @hooverst ????

OpenStudy (anonymous):

oh the generalized power rule is something like the chain rule. For instance, you sure know how to derivate \[x^2\] and you could also derive the following \(( x+3)^2\) by just multiplying the binomial out, but the generalized power rule works for higher orders two, for instance for \((x^2+4x+3)^{101}\) but we don't have to do such an example here, just make a substitution for the inside and derive it normally, then back substitute and multiply with the derived substitution.

OpenStudy (anonymous):

If you know about Leibniz Notation you can see how that works yourself.

OpenStudy (anonymous):

Let me show you how to do it anyway, you don't have to do all these steps every time you see such a function though, just be aware about that it is the generalized power rule or for all sorts of functions, as the chain rule \[y=(4x+3)^3 \ \text{let} \ u=4x+3 \rightarrow \frac{du}{dx}=4 \\y=u^3 \rightarrow \frac{dy}{du}=3u^2 \\ \frac{du}{dx}\cdot \frac{dy}{du}=\frac{dy}{dx}=12u^2 = 12(4x+3)^2\]

OpenStudy (anonymous):

ok got it.